PhotoLab charges $9.00 for processing a 24-exposure roll of film and $12.00 for processing a 36-exposure roll. After their class trip, Mr. Renkens' 8th grade class sent 22 rolls of film to be developed and paid $219 for processing.

A. Write a system of equations to describe the situation.

B. Find the number of rolls of 24-exposure and the number of 36-exposure film that were processed using either elimination or substitution. Make sure to show your work for full credit.

Answer by macston(5194)

X=number of 24-exposure rolls; Y=number of 35-exposure rolls
SECTION A: SYSTEM OF EQUATIONS
X+Y=22
$9.00X+$12.00Y=$219
SECTION B: SOLUTIONS
Method 1-Substitution
From equation1 above
X=22-Y Substitute for X in second equation.
$9.00X+$12.00Y=$219
$9.00(22-Y)+$12.00Y=$219
$198-$9.00Y+$12.00Y=$219 Subtract $198 from each side.
$3.00Y=$21 Divide each side by $3.00
Y=7 ANSWER 1: There were 7 rolls of 36-exposure film.
X+Y=22
X+7=22 Subtract 7 from each side.
X=15 ANSWER 2: There were 15 rolls of 24-exposure film.
CHECK:
$9.00X+$12.00Y=$219
$9.00(15)+$12.00(7)=$219
$135+$84=$219
$219=$219
Method 2-Elimination
X+Y=22
$9.00X+$12.00Y=$219 Multiply first equation by $9.00 and subtract
$9.00X+$9.00Y=$198
$3.00Y=$21 Divide each side by $3.00
Y=7 ANSWER 1: There were 7 rolls of 36-exposure film.
Continue as in Method 1